# Stability of vehicular admission control schemes in urban traffic networks under modelling uncertainty

### v1

@@ Abstract
Urban transportation networks face significant challenges due to traffic congestion, leading to adverse environmental and socioeconomic impacts. Vehicular admission control (VAC) strategies have emerged as a promising solution to alleviate congestion. By leveraging information and communication technologies, VAC strategies regulate vehicle entry into the network to optimize different traffic metrics of interest over space and time. Despite the significant development of VAC strategies, their stability at the presence of modelling uncertainty remains under-explored. This paper investigates the stability properties of a class of decentralized VAC schemes under modelling uncertainty. Specifically, we consider large-scale, heterogeneous urban traffic networks characterised by nonlinear dynamics and concave macroscopic fundamental diagrams with bounded uncertainty between flow, density, and speed. In this context, we examine a broad class of decentralized VAC dynamics, described by general nonlinear forms. Using passivity theory, we derive scalable, locally verifiable conditions on the design of VAC schemes, that enable stability guarantees in the presence of modelling uncertainty. Several examples are presented to illustrate the applicability of the proposed design framework. Our analytical results are validated through numerical simulations on a 6-region system, demonstrating their effectiveness and practical relevance.
@@ I Introduction
Traffic congestion remains one of the most critical challenges in urban transportation networks, leading to severe societal, economic, and environmental consequences. As urbanization accelerates and vehicle ownership rises, traffic demand increasingly exceeds available infrastructure capacity, particularly during peak hours, causing congestion. This imbalance can result in excessive delays and, in extreme cases, gridlock, where vehicular movement comes to a complete standstill[causes2]. To address these issues, researchers have extensively explored congestion mitigation strategies, ranging from traffic control measures, such as traffic signal optimization and variable speed limits, to demand-side interventions that influence traveller decisions on route choice and departure timing. Despite significant advancements, modelling and efficiently managing traffic in multi-regional urban networks remain challenging problems due to the dynamic interactions between different network regions, the unpredictability of human behaviour, and the limitations of existing infrastructure.
@@
Research on traffic flow modelling has been an active field since the first studies that identified the fundamental relations in 1934 and 1935[Greenshields1934THEPM, Greenshields1935ASO]. Early developments led to the emergence of both microscopic and macroscopic traffic models in the 1950s[Reuschel1950, Chandler1958, Leslie1961]. Microscopic models focus on the behaviour of individual vehicles, capturing interactions such as car-following and lane-changing dynamics. In contrast, macroscopic models describe traffic at an aggregated level, such as the road link level, treating vehicles as a continuous flow and using accumulated quantities like vehicle density, flow, and speed to represent the system’s state. Link-level macroscopic models are often preferred over microscopic models for large-scale real-time traffic control applications due to their reduced complexity[Orosz2010TrafficJD, WageningenKessels2015GenealogyOT]. However, their representational effectiveness is frequently undermined by significant uncertainties arising from complex vehicle interactions in urban environments.
@@
A key advancement was the development of the macroscopic fundamental diagram (MFD), which provides a low-scatter, unimodal relationship between traffic flow and vehicle accumulation in homogeneous traffic networks. The concept was first suggested by empirical observations in[Godfrey1969], which showed that every traffic network is characterized by a unique minimum average trip speed dictated by its physical structure. This laid the foundation for subsequent studies[Herman1979ATA, DAGANZO200749, GEROLIMINIS2008759, Helbing2009], which confirmed that urban traffic could be effectively described using MFDs. The MFD framework enables the partitioning of large-scale urban networks into homogeneous regions, where each region exhibits its own distinct MFD, capturing the relationship between flow and vehicle accumulation[DAGANZO200749]. This regional partitioning enables the development of effective real-time traffic control strategies[RAMEZANI20151, ZHANG20131, DAGANZO200749, geroliminis2013], as MFD-based approaches are robust to demand fluctuations and provide an aggregated yet effective representation of network dynamics.
@@
Traffic control aims to optimize vehicular flows across the network[papageorgioureview], employing strategies such as route guidance to redistribute traffic[geroleminisisik, MENELAOU2017, Lei2020], and gating and perimeter control (PC) to regulate inflows in protected regions[papageorgiou1990dynamic, papageorgioureview, ABOUDOLAS2013265, HADDAD2012, HADDAD2014, KOUVELAS2017, LI2020311, Lei2020, NING2023]. However, while protected regions remain congestion-free, other areas of the network, including the perimeter, experience exacerbated congestion[kouvelas2018hierarchical]. To address such problems Model Predictive Control (MPC) approaches have been utilized[mpc, geroliminis2013, SIRMATEL2021], which typically combine one or more of these strategies in a hierarchical manner to improve traffic conditions[Zhou2016, FU2017, Boufous2020]. While offering promising results, MPC schemes often overlook stability and uncertainty considerations and are computationally intensive.
@@
Despite the plethora of traffic control strategies and modern enabling technologies, congestion continues to be a persistent challenge[spectrum2019]. This is primarily driven by factors such as demand exceeding infrastructure capacity, inherent uncertainties in traffic modelling, and the occurrence of disruptive events. Vehicular admission control (VAC) strategies have emerged as promising solutions with the potential to significantly reduce or even eliminate congestion. By harnessing information and communication technologies, VAC regulates vehicle entry into the traffic network, such as by instructing vehicles when to depart from their origin, to optimize key traffic metrics [^1] While other works characterize vehicle departure time selection as a demand management strategy, we consider it as a VAC strategy. The reason is to emphasize that the objective is not to modify demand itself, but rather to regulate its entry into the system.. This ensures that waiting vehicles remain outside the network, preventing excessive accumulation, maintaining smooth traffic flow, and allowing users to fully utilize their time at their origin before departure[MENELAOU2017]. Although VAC may seem similar to PC, there is significant difference between these two control approaches. In particular, PC regulates the flows exchanged between a protected region and its neighbours, while VAC does not. Instead, VAC controls the spatio-temporal distribution of inflow in different regions from the outside world, but not the inter-regional flows.
@@
Early research on VAC focused on optimizing individual vehicle departure times and routes in a single region[MENELAOU2017, menelaouTRR, menelaouTIV]. Leveraging road reservations to predict future network conditions, [MENELAOU2017] considered the minimization of travel times on a per-vehicle basis, while ensuring that the regional density remained below the critical density, where maximum outflow is observed. Expanding on this approach,[menelaouTRR] introduced a composite objective to promote regional homogeneity, while[menelaouTIV] enhanced the model by predicting link speeds instead of assuming free-flow conditions. To accommodate large networks with multiple regions, several studies have adopted a macroscopic VAC framework that integrates regional dynamics[MENELAOU2022, MENELAOU2023, RAMP2024]. The works in[MENELAOU2022, MENELAOU2023] investigated triangular and nonlinear MFDs within MPC frameworks, developing tailored optimization algorithms to solve the resulting nonconvex problems, but without considering uncertainty or stability.
@@
Developing VAC strategies that account for uncertainty and guarantee stability remains a significant challenge. The critical role of stability was emphasized in[RAMP2024], which demonstrated that stability may be compromised when a region is operated near the critical density point, where even minor disturbances, such as small increases in demand or flow irregularities, may lead to congestion or even gridlock. Hence, an ongoing challenge is the identification of suitable conditions on VAC dynamics that ensure traffic network stability. Moreover, such conditions are important to be scalable, enabling VAC designs that require no recalibration when network dynamics or topology change, and distributed, minimizing computational and communication demands while avoiding a single point of failure. Another significant challenge arises from modelling uncertainty in MFDs, which stems from factors such as traffic heterogeneity, poor data quality, erratic driver behaviour, and unpredictable events like accidents and changing weather conditions. These challenges underscore the need to investigate the design and stability properties of VAC schemes in large-scale urban transportation networks where regional MFDs are uncertain.
@@
Contributions. This paper investigates the behaviour of traffic networks with heterogeneous regional, nonlinear, MFD dynamics in the presence of modelling uncertainty. The considered system is coupled with a broad class of nonlinear VAC dynamics. Using passivity theory, we derive conditions on the VAC dynamics that guarantee asymptotic stability. These conditions are distributed, relying only on locally available information, locally verifiable, and applicable to arbitrary connected network configurations, ensuring scalability. The applicability of the proposed approach is demonstrated with five VAC scheme examples that satisfy the developed conditions. Numerical simulations on a 6-region system highlight the effectiveness and practicality of the approach by showing network stability under suitable VAC schemes, even in the presence of modelling uncertainty and user non-adherent behaviour.
@@
Notation. The set of real numbers is given by $\mathbb{R}$, and the set of real non-negative numbers by $\mathbb{R}_{+}$. We denote vectors by bold small letters with the $i^{th}$ component of a vector $\mathbf{y}$ given by $y_{i}$. The set $\mathbb{R}^{n}$ contains all $n$-dimensional real vectors and the set $\mathbb{S}^{n}$ contains all the real, symmetric matrices of $n\times n$ dimension. The $n\times 1$ vector with all elements equal to 0 is given by $\mathbf{0}\in\mathbb{R}^{n}$. Capital calligraphic letters denote sets. The operation of removing an element $i$ from a set $\mathcal{K}$ is denoted by $\mathcal{K}\setminus\{i\}$. For $a,b\in\mathbb{R}$ and $a\leq b$, the min-max operator is given by $x]_{a}^{b}=\max(\min(x,b),a)$. A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is called positive definite if $f(0)=0$ and $f(x)>0$ for all $x\in\mathbb{R}^{n}$ elsewhere, and it is negative definite if $f(0)=0$ and $f(x)<0$ for every $x\neq 0$. The equilibrium of a system with state $x$ is denoted by $x^{*}$. We denote the Laplace transform of a function $q(t)$ by $\acute{q}=\int_{0}^{\infty}e^{-st}q(t)\,dt$. Units of variables are stated at first mention for clarity and conciseness.
@@ II-A Model description
The regions of the network are considered to be homogeneous in the sense that the aggregated traffic behaviour of each region ii [^2] Regions are heterogeneous among them since they can have different dynamics., is captured by a nonlinear macroscopic fundamental diagram (MFD)[DAGANZO200749, GEROLIMINIS2008759, KNOOP2015], see Fig. 2, that describes the total inter-regional flow, $g_{i}(\rho_{i})\;[{\text{veh}}/{\text{h}}]$ of the region ($\rho_{i}\;[{\text{veh}}/{\text{km}}]$ denotes vehicle density) as follows
@@
1(i) ensures that $d_{i}:[0,\rho_{i}^{J}]\to\mathbb{R}$ is bounded and continuous. This is to ensure that the resulting inter-regional flow values will capture real observations since MFDs describe a uniform, low-scatter relationship that is not affected by the change of demand[Godfrey1969, DAGANZO200749, GEROLIMINIS2008759]. 1(ii) ensures that the values of the uncertainty term will not result to negative inter-regional flow values, violating the physical behaviour that characterises MFD relationships. 1(iii) is needed to describe the behaviour of the total inter-regional flow, $g_{i}(\rho_{i})$, namely that even in a perturbed state, the MFD still shows two distinct flow regimes (free-flow, and congested flow) with a peak that can occur at a density that can be different from $\rho_{i}^{C,f}$[Godfrey1969, DAGANZO200749, GEROLIMINIS2008759]. In practical settings, and despite flow fluctuations or congestion, it has been demonstrated that the peaks of $g_{i}(\rho_{i})$ and $f_{i}(\rho_{i})$ occur at densities that are close[GEROLIMINIS2008759, KEYVANEKBATANI20121393, AMIRGHOLY2017261]. Assumptions 2(i-ii) ensure that the nonlinear function $f_{i}(\rho_{i}(t))$ yields values similar to real flow measurements and result in $f_{i}(\rho_{i}(t))$ accurately describing the two traffic conditions that characterize the total inter-regional flow of a region $i$, $g_{i}(\rho_{i}(t))$, see Fig. 2. For a region $i$, these traffic conditions are defined by the density value $\rho_{i}^{C,g}$ that corresponds to the maximum value of $g_{i}(\rho_{i})$, i.e., $g^{C}_{i}$. If $\rho_{i}(t)\leq\rho_{i}^{C,g}$ the region operates at free-flow conditions, while if $\rho_{i}(t)>\rho_{i}^{C,g}$ the region operates at congestion, with the behaviour produced by $g_{i}(\rho_{i})$ for both conditions consistent with real MFD observations[GEROLIMINIS2008759].
@@
Figure 2: General form of region $i$ total inter-regional traffic flow $g_{i}(\rho_{i})$ [veh/h], approximated by a nonlinear macroscopic fundamental diagram (MFD) area, due to flow variation and uncertainties[MENELAOU2023].
@@ III Vehicular admission control schemes design
We aim to design a class of decentralized VAC schemes that only utilize the in-region density, $\rho_{i}(t)$, since it is a direct measure for congestion that can be reliably estimated, as shown in[WANG8283631, DARWISH2015337]. This is done by considering the 1 admitted vehicular demand, $u_{i}(t)$, as the output of the following independent VAC dynamics
@@ III-B Passive admission control schemes
In this section, we propose suitable passivity properties, inspired from[WEN1266772, KASIS7776980], for the VAC dynamics, (10), to facilitate the analysis of their feedback interconnection with 1.
@@
The passivity properties of the control dynamics, (10), can be verified easily for general linear systems (see[KHALIL, KYP Lemma], transfer function positive realness) and for static nonlinearities. Explicit examples are presented in Section V. Additionally, it is noted that 3 allows the inclusion of a large class of control dynamics that satisfy the local passivity conditions detailed in 2.

### v2

@@ Abstract
Urban transportation networks face significant challenges due to traffic congestion, leading to adverse environmental and socioeconomic impacts. Vehicular admission control (VAC) strategies have emerged as a promising solution to alleviate congestion. By leveraging information and communication technologies, VAC strategies regulate vehicle entry into the network to optimize different traffic metrics of interest over space and time. Despite the significant development of VAC strategies, their stability at the presence of modelling uncertainty remains under-explored. This paper investigates the stability properties of a class of decentralized VAC schemes under modelling uncertainty. Specifically, we consider large-scale, heterogeneous urban traffic networks characterised by nonlinear dynamics and concave macroscopic fundamental diagrams with bounded uncertainty between flow, density, and speed. In this context, we examine a broad class of decentralized VAC dynamics, described by general nonlinear forms. Using passivity theory, we derive scalable, locally verifiable conditions on the design of VAC schemes, that enable stability guarantees in the presence of modelling uncertainty. Several examples are presented to illustrate the applicability of the proposed design framework. Our analytical results are validated through numerical simulations on a 6 and a 20-region system, demonstrating their effectiveness and practical relevance.
@@ I Introduction
Traffic congestion remains one of the most critical challenges in urban transportation networks, leading to severe societal, economic, and environmental consequences. As urbanization accelerates and vehicle ownership rises, traffic demand increasingly exceeds available infrastructure capacity, particularly during peak hours, causing congestion. This imbalance can result in excessive delays and, in extreme cases, gridlock, where vehicular movement comes to a complete standstill[3]. To address these issues, researchers have extensively explored congestion mitigation strategies, ranging from traffic control measures, such as traffic signal optimization and variable speed limits, to demand-side interventions that influence traveller decisions on route choice and departure timing. Despite significant advancements, modelling and efficiently managing traffic in multi-regional urban networks remain challenging problems due to the dynamic interactions between different network regions, the unpredictability of human behaviour, and the limitations of existing infrastructure.
@@
Research on traffic flow modelling has been an active field since the first studies that identified the fundamental relations in 1934 and 1935[16, 15]. Early developments led to the emergence of both microscopic and macroscopic traffic models in the 1950s[43, 44, 8]. Microscopic models focus on the behaviour of individual vehicles, capturing interactions such as car-following and lane-changing dynamics. In contrast, macroscopic models describe traffic at an aggregated level, such as the road link level, treating vehicles as a continuous flow and using accumulated quantities like vehicle density, flow, and speed to represent the system’s state. Link-level macroscopic models are often preferred over microscopic models for large-scale real-time traffic control applications due to their reduced complexity[38, 48]. However, their representational effectiveness is frequently undermined by significant uncertainties arising from complex vehicle interactions in urban environments.
@@
A key advancement was the development of the macroscopic fundamental diagram (MFD), which provides a low-scatter, unimodal relationship between traffic flow and vehicle accumulation in homogeneous traffic networks. The concept was first suggested by empirical observations in[12], which showed that every traffic network is characterized by a unique minimum average trip speed dictated by its physical structure. This laid the foundation for subsequent studies[20, 6, 10, 19], which confirmed that urban traffic could be effectively described using MFDs. The MFD framework enables the partitioning of large-scale urban networks into homogeneous regions, where each region exhibits its own distinct MFD, capturing the relationship between flow and vehicle accumulation[6]. This regional partitioning enables the development of effective real-time traffic control strategies[41, 51, 6, 11], as MFD-based approaches are robust to demand fluctuations and provide an aggregated yet effective representation of network dynamics.
@@
Traffic control aims to optimize vehicular flows across the network[39], employing strategies such as route guidance to redistribute traffic[46, 32, 29], and gating and perimeter control (PC) to regulate inflows in protected regions[40, 39, 1, 17, 18, 27, 30, 29, 37]. However, while protected regions remain congestion-free, other areas of the network, including the perimeter, experience exacerbated congestion[28]. To address such problems Model Predictive Control (MPC) approaches have been utilized[5, 11, 47], which typically combine one or more of these strategies in a hierarchical manner to improve traffic conditions[52, 9, 4]. While offering promising results, MPC schemes often overlook stability and uncertainty considerations and are computationally intensive.
@@
Despite the plethora of traffic control strategies and modern enabling technologies, congestion continues to be a persistent challenge[31]. This is primarily driven by factors such as demand exceeding infrastructure capacity, inherent uncertainties in traffic modelling, and the occurrence of disruptive events. Vehicular admission control (VAC) strategies have emerged as promising solutions with the potential to significantly reduce or even eliminate congestion. By harnessing information and communication technologies, VAC regulates vehicle entry into the traffic network, such as by instructing vehicles when to depart from their origin, to optimize key traffic metrics [^1] While other works characterize vehicle departure time selection as a demand management strategy, we consider it as a VAC strategy. The reason is to emphasize that the objective is not to modify demand itself, but rather to regulate its entry into the system.. This ensures that waiting vehicles remain outside the network, preventing excessive accumulation, maintaining smooth traffic flow, and allowing users to fully utilize their time at their origin before departure[32]. Although VAC may seem similar to PC, there is significant difference between these two control approaches. In particular, PC regulates the flows exchanged between a protected region and its neighbours, while VAC does not. Instead, VAC controls the spatio-temporal distribution of inflow in different regions from the outside world, but not the inter-regional flows.
@@
Early research on VAC focused on optimizing individual vehicle departure times and routes in a single region[32, 34, 33]. Leveraging road reservations to predict future network conditions, [32] considered the minimization of travel times on a per-vehicle basis, while ensuring that the regional density remained below the critical density, where maximum outflow is observed. Expanding on this approach,[34] introduced a composite objective to promote regional homogeneity, while[33] enhanced the model by predicting link speeds instead of assuming free-flow conditions. To accommodate large networks with multiple regions, several studies have adopted a macroscopic VAC framework that integrates regional dynamics[35, 36, 42]. The works in[35, 36] investigated triangular and nonlinear MFDs within MPC frameworks, developing tailored optimization algorithms to solve the resulting nonconvex problems, but without considering uncertainty or stability.
@@
Developing VAC strategies that account for uncertainty and guarantee stability remains a significant challenge. The critical role of stability was emphasized in[42], which demonstrated that stability may be compromised when a region is operated near the critical density point, where even minor disturbances, such as small increases in demand or flow irregularities, may lead to congestion or even gridlock. Hence, an ongoing challenge is the identification of suitable conditions on VAC dynamics that ensure traffic network stability. Moreover, such conditions are important to be scalable, enabling VAC designs that require no recalibration when network dynamics or topology change, and distributed, minimizing computational and communication demands while avoiding a single point of failure. Another significant challenge arises from modelling uncertainty in MFDs, which stems from factors such as traffic heterogeneity, poor data quality, erratic driver behaviour, and unpredictable events like accidents and changing weather conditions. These challenges underscore the need to investigate the design and stability properties of VAC schemes in large-scale urban transportation networks where regional MFDs are uncertain.
@@
Contributions. This paper investigates the behaviour of traffic networks with heterogeneous regional, nonlinear, MFD dynamics in the presence of modelling uncertainty. The considered system is coupled with a broad class of nonlinear VAC dynamics. Using passivity theory, we derive conditions on the VAC dynamics that guarantee asymptotic stability. These conditions are distributed, relying only on locally available information, locally verifiable, and applicable to arbitrary connected network configurations, ensuring scalability. The applicability of the proposed approach is demonstrated with five VAC scheme examples that satisfy the developed conditions. Numerical simulations on a 6 and a 20-region system highlight the effectiveness and practicality of the approach by showing network stability under suitable VAC schemes, even in the presence of modelling uncertainty and user non-adherent behaviour.
@@
Notation. The set of real numbers is given by $\mathbb{R}$, and the set of real non-negative numbers by $\mathbb{R}_{+}$. We denote vectors by bold small letters with the $i^{th}$ component of a vector $\mathbf{y}$ given by $y_{i}$. The set $\mathbb{R}^{n}$ contains all $n$-dimensional real vectors and the set $\mathbb{S}^{n}$ contains all the real, symmetric matrices of $n\times n$ dimension. The $n\times 1$ vector with all elements equal to 0 is given by $\mathbf{0}\in\mathbb{R}^{n}$. Capital calligraphic letters denote sets. The operation of removing an element $i$ from a set $\mathcal{K}$ is denoted by $\mathcal{K}\setminus\{i\}$. For $a,b\in\mathbb{R}$ and $a\leq b$, the min-max operator is given by $x]_{a}^{b}=\max(\min(x,b),a)$. A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is called positive definite if $f(0)=0$ and $f(x)>0$ for all $x\in\mathbb{R}^{n}$ elsewhere, and it is negative definite if $f(0)=0$ and $f(x)<0$ for every $x\neq 0$. The equilibrium of a system with state $x$ is denoted by $x^{*}$. The Laplace transform of a system with input $\rho(t)$ and output $u(t)$ is denoted by G(s). Units of variables are stated at first mention for clarity and conciseness.
@@ II-A Model description
The regions of the network are considered to be homogeneous in the sense that the aggregated traffic behaviour of each region ii [^2] Regions are heterogeneous among them since they can have different dynamics., is captured by a nonlinear macroscopic fundamental diagram (MFD)[6, 10, 26], see Fig. 2, that describes the total inter-regional flow, $g_{i}(\rho_{i})\;[{\text{veh}}/{\text{h}}]$ of the region ($\rho_{i}\;[{\text{veh}}/{\text{km}}]$ denotes vehicle density) as follows
@@
1(i) ensures that $d_{i}:[0,\rho_{i}^{J}]\to\mathbb{R}$ is bounded and continuous. This is to ensure that the resulting inter-regional flow values will capture real observations since MFDs describe a uniform, low-scatter relationship that is not affected by the change of demand[12, 6, 10]. 1(ii) ensures that the values of the uncertainty term will not result to negative inter-regional flow values, violating the physical behaviour that characterises MFD relationships. 1(iii) is needed to describe the behaviour of the total inter-regional flow, $g_{i}(\rho_{i})$, namely that even in a perturbed state, the MFD still shows two distinct flow regimes (free-flow, and congested flow) with a peak that can occur at a density that can be different from $\rho_{i}^{C,f}$[12, 6, 10]. In practical settings, and despite flow fluctuations or congestion, it has been demonstrated that the peaks of $g_{i}(\rho_{i})$ and $f_{i}(\rho_{i})$ occur at densities that are close[10, 24, 2]. Assumptions 2(i-ii) ensure that the nonlinear function $f_{i}(\rho_{i}(t))$ yields values similar to real flow measurements and result in $f_{i}(\rho_{i}(t))$ accurately describing the two traffic conditions that characterize the total inter-regional flow of a region $i$, $g_{i}(\rho_{i}(t))$, see Fig. 2. For a region $i$, these traffic conditions are defined by the density value $\rho_{i}^{C,g}$ that corresponds to the maximum value of $g_{i}(\rho_{i})$, i.e., $g^{C}_{i}$. If $\rho_{i}(t)\leq\rho_{i}^{C,g}$ the region operates at free-flow conditions, while if $\rho_{i}(t)>\rho_{i}^{C,g}$ the region operates at congestion, with the behaviour produced by $g_{i}(\rho_{i})$ for both conditions consistent with real MFD observations[10].
@@
Figure 2: General form of region $i$ total inter-regional traffic flow $g_{i}(\rho_{i})$ [veh/h], approximated by a nonlinear macroscopic fundamental diagram (MFD) area, due to flow variation and uncertainties[36].
@@ III Vehicular admission control schemes design
We aim to design a class of decentralized VAC schemes that only utilize the in-region density, $\rho_{i}(t)$, since it is a direct measure for congestion that can be reliably estimated, as shown in[49, 7]. This is done by considering the 1 admitted vehicular demand, $u_{i}(t)$, as the output of the following independent VAC dynamics
@@ III-B Passive admission control schemes
In this section, we propose suitable passivity properties, inspired from[50, 23], for the VAC dynamics, (10), to facilitate the analysis of their feedback interconnection with 1.
@@
The passivity properties of the control dynamics, (10), can be verified easily for general linear systems (see[25, KYP Lemma], transfer function positive realness) and for static nonlinearities. Explicit examples are presented in Section V. Additionally, it is noted that 3 allows the inclusion of a large class of control dynamics that satisfy the local passivity conditions detailed in 2.